OpenStudy (anonymous):

The first part of the 660 mile journey was by car at 30 mph. The 2nd part was by bus @ 60mph. If the total time was 14 hours, what was the distance traveled by bus?

OpenStudy (anonymous):

Let the time the car traveled be t. Then, using (distance)=(speed)*(time) for each part of the journey,\[660=30t+60(14-t)\]Solve for t. That's the time taken by the car. You need 'bus time', so subtract t from 14 hrs.

OpenStudy (anonymous):

\[t_{bus}=8h\]

OpenStudy (amistre64):

30x + 60y = 660; and x+y=14 (hours) x=(14-y) 30(14-y) +60y = 660 420 -30y +60y = 660 y(60-30) = 660-420 30y = 240 y= 240/30 = 24/3 = 8 The bus ride was 8 hrs long

OpenStudy (anonymous):

thanks!!!